A microwave oven of unknown resistance is connected to a 110V voltage source.calculate the oven’s resistance if the current needed has an intensity of 7,200mA
1
Expert's answer
2018-05-31T07:13:08-0400
Solution:
Ohm law R=U/I=110/7.2=15.28 Ohm